OpenStudy (anonymous):

Solving Equations 8+4m=9m-7 Helpp!

OpenStudy (anonymous):

post the options

OpenStudy (anonymous):

theres no options

OpenStudy (anonymous):

m=3

OpenStudy (shamil98):

8 + 4m = 9m - 7 you want to isolate m.

OpenStudy (gunnar):

subtract 4 on 9 = 5 then 8+7 = 15 15/5= 3

OpenStudy (shamil98):

8 + 4m = 9m -7 add 7 to both sides 15 + 4m = 9m subtract 4m to both sides 15 = 5m divide by 5 3 = m

OpenStudy (anonymous):

ok... m=3

OpenStudy (anonymous):

thanks!;P